OpenStudy (anonymous):

Find the solution of this system of equations. Seperate the x- and y-values with a comma. x=8+y -2x+6y=-20

pooja195 (pooja195):

Use substiotuion :)

pooja195 (pooja195):

\[\huge-2(8+y)+6y=-20\]

OpenStudy (z4k4r1y4):

solve the first equation for 'y' then substitute into second equation.

OpenStudy (anonymous):

I'm not good with this stuff at all lol

pooja195 (pooja195):

Distribute first \[\huge −2(8+y)\]

OpenStudy (anonymous):

-16 - 2y

pooja195 (pooja195):

good :) \[\huge-16-2y+6y\] combine like terms

OpenStudy (anonymous):

-16+4

pooja195 (pooja195):

no

OpenStudy (anonymous):

-16+4y*

pooja195 (pooja195):

good \[4y-16=-20\] isolate y

OpenStudy (anonymous):

I don't know how to do any of that haha, I'm just finishing a study guide I have so I can pass for the year :/

pooja195 (pooja195):

Well firt step to isolate would be to add the 16 to both sides :) -16+16=0 so it cancels -20+16=?

OpenStudy (anonymous):

-4

pooja195 (pooja195):

good now just divide! :) 4y=-4 4/4=1 so you are left with y -4/4=?

OpenStudy (anonymous):

-1

OpenStudy (anonymous):

so it's 7, -1?

pooja195 (pooja195):

yep :) good job!

pooja195 (pooja195):

\(\huge\color{orange}{Welcome~To~OpenStudy!!~:)}\) Hey there!!! Since you are new here, read this legendary tutorial for new OpenStudiers!! http://goo.gl/5pp1u0 For an visual tutorial look at this! Its amazing :o https://prezi.com/fs3hqdpcopic/an-unofficial-guide-to-openstudy/ Also consider reading the CoC and the T&C http://openstudy.com/code-of-conduct http://openstudy.com/terms-and-conditions

OpenStudy (anonymous):

Thanks!